# SIGMA: Selective Gated Mamba for Sequential Recommendation

### v3

@@ A. Complexity Analysis
In this section, we will analyze the time complexity of our proposed SIGMA framework. We denote training iteration as ${I}$, the sequence length as ${N}$, the embedding size as ${D}$ and the kernel size of Conv1d as $k$. Following the previous formulas, we analyze the time complexity of the key components: (i) Firstly, since our DS Gate consists of several linear layers and one Conv1d Layer, the computation cost of it can be calculated as $\mathcal{O}\left(3\times N\times D^{2}+N\times D\times k\right)$. (ii) Secondly, the time complexity of Mamba is known as $\mathcal{O}\left(N\times D\right)$ (Gu and Dao 2023). In the PF-Mamba, we add one more reverse direction and several linear layers, therefore, the total complexity can be calculated as $\mathcal{O}\left(3\times N\times D^{2}+N\times D\times(k+2)\right)$. (iii) Thirdly, for the FE-GRU, which consists of Conv1d and GRU cell, the computational complexity can be calculated by simply adding GRU and Conv1d together: $\mathcal{O}\left(N\times D^{2}+N\times D\times k\right)$. In conclusion, the whole time complexity of SIGMA is $\mathcal{O}\left(13\times N\times D^{2}+N\times D\times(2k+2)\right)$. Considering the extreme situation when dealing with quite long sequences ($\left(N\gg D\right)$), it can be further simplified to $\mathcal{O}\left(N\right)$, compared to the $\mathcal{O}\left({N}^{2}\right)$ complexity of transformer-based models (Liu et al. 2024a), showing its superiority in efficiency, which also support by the experimental results on ML-1M dataset.
@@ B. Dataset Information
We mainly evaluate our framework on five real-world datasets, i.e., Beauty, Sports, Games, Yelp and ML-1M, which are all large-scale public datasets that have been widely used as benchmarks in the next-item prediction task (Yang et al. 2024).
@@
In the experiments, we employ a leave-one-out method for splitting the datasets and arrange all the user interactions sequentially by time. Moreover, for each user and item, we construct an interaction sequence by simply sorting their interaction records based on timestamps and ratings. Considering the average length and total samples of each dataset vary, we filter users and items with less than five recorded interactions for ML-1M, Beauty, and Games, following the setting in original papers (Qin, Yang, and Zhong 2024). For Yelp and Sports, we also added upper bounds ($100$ for Yelp and $200$ for Sports).
@@ E. Baselines
(a) GRU4Rec (Vaswani et al. 2017): GRU4Rec utilizes Gated Recurrent Units (GRUs) to capture sequential dependencies within user interaction data. It is particularly effective for session-based recommendations, allowing the model to focus on the most recent and relevant user interactions to make accurate predictions. The model is known for its efficiency in handling session-based data, particularly in capturing user intent during a session (Keles, Wijewardena, and Hegde 2023). (b) BERT4Rec (Kang and McAuley 2018): BERT4Rec adapts the BERT (Bidirectional Encoder Representations from Transformers) architecture for personalized recommendations. Unlike traditional sequential models, BERT4Rec considers both past and future contexts of user behavior by employing a bidirectional self-attention mechanism, which enhances the representation of user interaction sequences. This allows the model to predict items in a more context-aware manner. (c) SASRec (Qin, Yang, and Zhong 2024): SASRec applies a multi-head self-attention mechanism to capture both long-term and short-term user preferences from interaction sequences. It constructs user representations by focusing on relevant parts of the interaction history, thereby improving the quality of recommendations, especially in scenarios where user preferences vary over time. (d) LinRec (de Souza Pereira Moreira et al. 2021): LinRec simplifies the computational complexity of the traditional transformer models by modifying the dot product in the attention mechanism. This reduction in complexity makes LinRec particularly suitable for large-scale recommendation tasks where efficiency is crucial, without significantly sacrificing performance. (e) FEARec (Kim et al. 2019a): FEARec enhances traditional attention mechanisms by incorporating information from the frequency domain. This hybrid approach allows the model to better capture periodic patterns and long-range dependencies in user interaction sequences, leading to more powerful and accurate recommendations. (f) Mamba4Rec (Jannach and Ludewig 2017): Mamba4Rec leverages Selective State Space Models (SSMs) to address the effectiveness-efficiency trade-off in sequential recommendation. By using SSMs, Mamba4Rec efficiently handles long behavior sequences, capturing complex dependencies while maintaining low computational costs. It outperforms traditional self-attention mechanisms, especially in scenarios involving long user interaction histories. (g) EchoMamba4Rec (12): Building on Mamba4Rec, EchoMamba4Rec introduces a frequency domain filter to remove noise and enhance the signal in sequential data. This bi-directional model processes sequences both forward and backward, providing a more comprehensive understanding of user behavior. The Fourier Transform-based filtering further improves the model’s robustness and accuracy in predicting user preferences.
@@ G. Grouped Users Analysis
Figure 7: Grouped Users Analysis on ML-1M, Yelp and Games.
@@ I. Guideline for reproducibility
In this section, we will provide detailed guidelines for reproducing our results. In the model file we released, we have baseline and baseline_config folders, which respectively contain almost all the baseline methods (others can be easily found in recbole’s original models) and corresponding configurations mentioned in the overall experiment; datasets folder storing the chosen datasets (Beauty, Sports, Games, Yelp and ML-1M) in atomic files to fit the recbole framework (Wang, He, and Zhu 2024); model file containing the gated_mamba.py, which is the main structure of our SIGMA. To reproduce our proposed SIGMA framework, you can directly run the run.py file with the proper command. It is noteworthy that the environment and version of mamba_ssm are quite important in reproducing the experimental results, so please check your environment referencing the requirements.yaml file. Specifically, you are recommended to run run.py for successful reproduction, which contains the calls to gated_mamba.py. For the relevant parameters to perform the experiments in hyperparameter analysis, you can directly set it in config.yaml with other dataset settings and model settings. We also attach the RecMamba.ipynb to show the raw training procedure in Colab.

### v4

@@ A. Complexity Analysis
In this section, we will analyze the time complexity of our proposed SIGMA framework. We denote training iteration as ${I}$, the sequence length as ${N}$, the embedding size as ${D}$ and the kernel size of Conv1d as $k$. Following the previous formulas, we analyze the time complexity of the key components: (i) Firstly, since our DS Gate consists of several linear layers and one Conv1d Layer, the computation cost of it can be calculated as $\mathcal{O}\left(3\times N\times D^{2}+N\times D\times k\right)$. (ii) Secondly, the time complexity of Mamba is known as $\mathcal{O}\left(N\times D\right)$ (Gu and Dao 2023). In the PF-Mamba, we add one more reverse direction and several linear layers, therefore, the total complexity can be calculated as $\mathcal{O}\left(3\times N\times D^{2}+N\times D\times(k+2)\right)$. (iii) Thirdly, for the FE-GRU, which consists of Conv1d and GRU cell, the computational complexity can be calculated by simply adding GRU and Conv1d together: $\mathcal{O}\left(N\times D^{2}+N\times D\times k\right)$. In conclusion, the whole time complexity of SIGMA is $\mathcal{O}\left(13\times N\times D^{2}+N\times D\times(2k+2)\right)$. Considering the extreme situation when dealing with quite long sequences ($\left(N\gg D\right)$), it can be further simplified to $\mathcal{O}\left(N\right)$, compared to the $\mathcal{O}\left({N}^{2}\right)$ complexity of transformer-based models (Keles, Wijewardena, and Hegde 2023), showing its superiority in efficiency, which also support by the experimental results on ML-1M dataset.
@@ B. Dataset Information
We mainly evaluate our framework on five real-world datasets, i.e., Beauty, Sports, Games, Yelp and ML-1M, which are all large-scale public datasets that have been widely used as benchmarks in the next-item prediction task (Wang et al. 2019).
@@
In the experiments, we employ a leave-one-out method for splitting the datasets and arrange all the user interactions sequentially by time. Moreover, for each user and item, we construct an interaction sequence by simply sorting their interaction records based on timestamps and ratings. Considering the average length and total samples of each dataset vary, we filter users and items with less than five recorded interactions for ML-1M, Beauty, and Games, following the setting in original papers (Wang, He, and Zhu 2024). For Yelp and Sports, we also added upper bounds ($100$ for Yelp and $200$ for Sports).
@@ E. Baselines
(a) GRU4Rec (Jannach and Ludewig 2017): GRU4Rec utilizes Gated Recurrent Units (GRUs) to capture sequential dependencies within user interaction data. It is particularly effective for session-based recommendations, allowing the model to focus on the most recent and relevant user interactions to make accurate predictions. The model is known for its efficiency in handling session-based data, particularly in capturing user intent during a session (Li et al. 2017). (b) BERT4Rec (Sun et al. 2019): BERT4Rec adapts the BERT (Bidirectional Encoder Representations from Transformers) architecture for personalized recommendations. Unlike traditional sequential models, BERT4Rec considers both past and future contexts of user behavior by employing a bidirectional self-attention mechanism, which enhances the representation of user interaction sequences. This allows the model to predict items in a more context-aware manner. (c) SASRec (Kang and McAuley 2018): SASRec applies a multi-head self-attention mechanism to capture both long-term and short-term user preferences from interaction sequences. It constructs user representations by focusing on relevant parts of the interaction history, thereby improving the quality of recommendations, especially in scenarios where user preferences vary over time. (d) LinRec (Liu et al. 2023a): LinRec simplifies the computational complexity of the traditional transformer models by modifying the dot product in the attention mechanism. This reduction in complexity makes LinRec particularly suitable for large-scale recommendation tasks where efficiency is crucial, without significantly sacrificing performance. (e) FEARec (Du et al. 2023): FEARec enhances traditional attention mechanisms by incorporating information from the frequency domain. This hybrid approach allows the model to better capture periodic patterns and long-range dependencies in user interaction sequences, leading to more powerful and accurate recommendations. (f) Mamba4Rec (Liu et al. 2024a): Mamba4Rec leverages Selective State Space Models (SSMs) to address the effectiveness-efficiency trade-off in sequential recommendation. By using SSMs, Mamba4Rec efficiently handles long behavior sequences, capturing complex dependencies while maintaining low computational costs. It outperforms traditional self-attention mechanisms, especially in scenarios involving long user interaction histories. (g) EchoMamba4Rec (Wang, He, and Zhu 2024): Building on Mamba4Rec, EchoMamba4Rec introduces a frequency domain filter to remove noise and enhance the signal in sequential data. This bi-directional model processes sequences both forward and backward, providing a more comprehensive understanding of user behavior. The Fourier Transform-based filtering further improves the model’s robustness and accuracy in predicting user preferences.
@@ F. Efficiency Comparison
Figure 7: Grouped Users Analysis on ML-1M, Yelp and Games.
@@ I. Guideline for reproducibility
In this section, we will provide detailed guidelines for reproducing our results. In the model file we released, we have baseline and baseline_config folders, which respectively contain almost all the baseline methods (others can be easily found in recbole’s original models) and corresponding configurations mentioned in the overall experiment; datasets folder storing the chosen datasets (Beauty, Sports, Games, Yelp and ML-1M) in atomic files to fit the recbole framework (Zhao et al. 2021); model file containing the gated_mamba.py, which is the main structure of our SIGMA. To reproduce our proposed SIGMA framework, you can directly run the run.py file with the proper command. It is noteworthy that the environment and version of mamba_ssm are quite important in reproducing the experimental results, so please check your environment referencing the requirements.yaml file. Specifically, you are recommended to run run.py for successful reproduction, which contains the calls to gated_mamba.py. For the relevant parameters to perform the experiments in hyperparameter analysis, you can directly set it in config.yaml with other dataset settings and model settings. We also attach the RecMamba.ipynb to show the raw training procedure in Colab.